A gas occupies 195 cm3 at 43°C. Find its volume at 75°C



1
Expert's answer
2021-05-28T00:57:03-0400

"V_1=195cm^3 , T_1=43^oC=316K"

"T_2=75^oC=348K"

"V_2=?"

We will use the relation

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

"\frac{195}{316}=\frac{V_2}{348}"

"\boxed{V_2=214.74 cm^3}answer"